# Rethinking Spectral Graph Neural Networks with Spatially Adaptive Filtering

### 17 Jan 2024 v1

@@ C.3 Practical Significance of eigendecomposition
Eigendecomposition, or spectral decomposition, breaks down a matrix into its eigenvalues and eigenvectors, offering insights into matrix properties, especially for the graph Laplacian. Even with its computational demands, this technique has attracted surging interest in the graph learning community due to its theoretical richness and can be practically expedited for larger graphs using Sparse Generalized Eigenvalue algorithms (Cai et al., 2021). Recent advancements underscore its value in various applications such as graph positional encoding (Belkin & Niyogi, 2003; Wang et al., 2022; Dwivedi et al., 2020; Lim et al., 2022), spectral graph convolution (Bo et al., 2023; Liao et al., 2019), graph domain adaptation (You et al., 2022), and graph robustness (Chang et al., 2021). For example, Laplacian eigenvectors have be widely used in identifying nodes’ global position in the graph (Wang et al., 2022), particularly in recent popular graph transformers (Kreuzer et al., 2021; Kim et al., 2022; Rampášek et al., 2022), enhancing their expressiveness. Innovations like SignNet and BasisNet (Lim et al., 2022) have further optimized the processing of these eigenvectors. When exploring the expressive power of GNN models, Specformer (Bo et al., 2023) employs eigendecomposition for learning set-to-set graph filters, while FE-GNN (Sun et al., 2023) taps into singular value decomposition (SVD) for graph feature expansion. In line with these developments, our method, SAF, utilizes eigendecomposition to explicitly create a new graph, enabling efficient non-local aggregation to capture long-range dependencies and address graph heterophily.
Appendix D Dataset Information
@@
We conduct experiments on 13 real-world datasets from various domains. The detailed statistics are summarized in Table 3. Alongside common data attributes, we also provide the longest geodesic (shortest-path) distance between graph nodes for better illustrating the non-local property we investigate in Figure 1 and Figure 5. Moreover, we employ three metrics - edge homophily (Zhu et al., 2020a) $\mathcal{H}$, class homophily (Lim et al., 2021) $\mathcal{H}_{\text{class}}$, and adjusted homophily (Platonov et al., 2023) $\mathcal{H}_{\text{adjusted}}$ to assess the graph’s homophily ratio, which ranges from 0 (high heterophily) to 1 (high homophily). While the first is a commonly used index, the latter two, considering class variability and potential imbalance, have been recently introduced to provide a more accurate estimation. For our main text analysis regarding the adapted new graph, we primarily rely on the edge homophily metric, defined as =|$\mathcal{H}=|\{(v_{i},v_{j})|(v_{i},v_{j})\in\mathcal{E}\wedge\mathbf{y}_{i}=\mathbf{y}_{j}\}|/|\mathcal{E}|$, given its simplicity and wide usage. In certain compact sections of this paper, we use four-letter abbreviations for dataset names. As these datasets are open benchmarks, we have downloaded and utilized them from the respective web links provided below.
Appendix E Experimental Details
@@
In this section, we provide experimental details for reproducibility. As He et al. (2022) have made a comprehensive evaluation and share the same experimental protocol with us, we directly leverage their results for models: MLP, GCN, ARMA, APPNP, GCNII, PDE-GCN, GPR-GNN, BernNet, and ChebNetII on datasets including Chameleon, Squirrel, Texas, Cornell, Actor, Cora, Citeseer, and Pubmed. Other experiments are performed on a machine equipped with an NVIDIA GeForce RTX 3090 (24GB) and an Intel(R) Xeon(R) Gold 5218R CPU @ 2.10GHz (20 cores).
E.1 Our Implementations
The main codes for our proposed SAF can be founded in the supplementary material.
E.2 Baseline Implementations
@@
In our experiments, we leverage the Pytorch Geometric library [^2] https://github.com/pyg-team/pytorch_geometric implementations for GCN and APPNP. For MLP, we include a sequence of linear layers, each of which is followed by batch normalization, ReLU activation, and dropout. The number of MLP layers are tuned from 1 to 5. For the remaining baselines, we resort to the officially released code, accessible via the provided URLs. Besides, for the work (Li et al., 2022), only the most effective model variant, GloGNN++, is included in our experiments.
E.3 Hyper-parameters Setting
@@
Appendix F More Experiments
F.1 Empirical Evidence of the Adapted New Graph
Additional empirical evidence related to the new graph’s non-locality and signed edge weights can be found in Figure 5 and Figure 6, respectively.
Figure 5: Additional distributions of original hop distance between adjacent nodes in the adapted new graph, where the number of bins is set as five for better histgram visualization.
Figure 6: Additional visualizations for illustrating the newfound graph’s ability in discerning label consistency. Blue, yellow, and green bars (sharing the same meanings as in the main text) respectively signifies the original graph, positive edges of new graph, and negative edges of new graph.
F.2 New Benchmarks for Graph Heterophily.
For a more extensive evaluation across various domains, we also test SAF on the recently introduced datasets, including Minesweeper, Tolokers, Amazon-ratings, Roman-empire, and Twitch-DE. In this context, we draw comparisons solely with MLP, GCN, APPNP, along with six methods that have previously shown promising results in prior tasks, namely NodeFormer, GloGNN++, GPR-GNN, BernNet, ChebNetII, JacobiConv, and FE-GNN. Table 4 lists the average classification accuracies, obtained over 10 random splits as provided by (Platonov et al., 2023), with a distribution of 50%/25%/25% for training/validation/testing. To maintain consistency with this setting, we randomly divide the Twitch-DE dataset (Lim et al., 2021) ourselves. In summary, SAF achieves significant performance gains of 12.84% and 4.10% on Minesweeper and Tolokers, respectively. Meanwhile, SAF maintains competitiveness on the other three datasets, validating its reliability and effectiveness.
Table 4: Experimental results on new heterophilic graph datasets.
Method Minesweeper Tolokers Amazon-ratings Roman-empire Twitch-DE
MLP 50.61$\pm$0.87 74.58$\pm$0.69 45.50$\pm$0.38 66.11$\pm$0.33 68.50$\pm$0.87
GCN 72.25$\pm$0.60 76.56$\pm$0.85 48.06$\pm$0.39 53.49$\pm$0.33 73.83$\pm$0.69
APPNP 68.48$\pm$1.20 74.13$\pm$0.62 48.12$\pm$0.37 72.99$\pm$0.46 72.81$\pm$0.66
NodeFormer 89.89$\pm$0.46 80.31$\pm$0.75 43.67$\pm$1.54 73.59$\pm$0.60 69.66$\pm$0.83
GloGNN++ 72.59$\pm$1.54 79.01$\pm$0.84 50.03$\pm$0.29 66.10$\pm$0.26 73.15$\pm$0.59
GPR-GNN 89.76$\pm$0.53 75.82$\pm$0.50 49.06$\pm$0.25 73.19$\pm$0.24 73.48$\pm$0.81
ChebNetII 83.62$\pm$1.51 78.95$\pm$0.49 49.76$\pm$0.36 74.52$\pm$0.54 74.17$\pm$0.68
JacobiConv 89.88$\pm$0.33 77.24$\pm$0.39 43.89$\pm$0.28 74.30$\pm$0.50 74.11$\pm$0.75
FE-GNN 84.68$\pm$0.36 79.31$\pm$0.59 49.46$\pm$0.29 74.50$\pm$0.30 73.60$\pm$0.60
BernNet 77.75$\pm$0.61 75.35$\pm$0.63 49.84$\pm$0.52 74.56$\pm$0.74 73.87$\pm$0.63
SAF 90.54$\pm$0.30 79.38$\pm$0.58 50.49$\pm$0.28 74.87$\pm$0.22 74.65$\pm$0.65
SAF-$\epsilon$ 90.59$\pm$0.35 79.45$\pm$0.60 50.36$\pm$0.15 74.89$\pm$0.29 74.55$\pm$0.68
Improv. 12.84% 4.10% 0.65% 0.33% 0.78%
F.3 Ablation Study
Our SAF framework primarily comprises three modules: “Non-negative Spectral Filtering” (Spec.), “Non-local Spatial Aggregation” (Spat.), and “Node-wise Prediction Amalgamation” (Amal.). While the ablation study of the Spec. module is thoroughly analyzed in Section 6 of the main text, this appendix section extends this analysis by offering additional visualization results on more datasets, as depicted in Figure 7.
Focusing on the other modules, detailed results are presented in Table 5. These findings consistently show that excluding any module typically leads to a noticeable drop in performance, affirming the critical role each plays in our SAF framework. Notably, for the Amal. module, certain datasets, such as Cham. and Squi., demonstrate only a slight reduction in performance. This observation aligns with our observation that their optimal attention values are close to an even split, as suggested in Figures 2(b) and 7(a).
Figure 7: Ablation study of SAF w.r.t. Spectral Filtering on the additional seven datasets.
Table 5: Ablation study of SAF framework on 13 datasets regarding “Non-negative Spectral Filtering” (Spec.), “Non-local Spatial Aggregation” (Spat.) and “Node-wise Prediction Amalgamation” (Amal.) modules. For SAF w/o Amal., we remove the attention modules and blend predictions from different domains equally. As the SAF w/o Spat. configuration is equivalent to BernNet model, the corresponding results are posted directly.
Our Variant Cham. Squi. Texas Corn. Actor Cora Cite. Pubm. Mine. Tolo. Amaz. Roma. Twit.
SAF-$\epsilon$ 74.84 64.00 94.75 93.28 42.98 89.87 81.45 91.52 90.59 79.45 50.36 74.89 74.55
SAF 75.30 63.63 94.10 92.95 42.93 89.80 80.61 91.49 90.54 79.38 50.49 74.87 74.65
SAF w/o Amal. 75.01 62.62 89.18 86.07 41.53 88.80 80.37 91.24 89.45 76.18 49.98 74.25 73.05
SAF w/o Spec. 73.55 55.70 90.49 88.20 41.06 88.03 78.87 90.12 89.45 78.23 49.13 71.85 73.54
SAF w/o Spat. 68.53 51.39 92.62 92.13 41.71 88.51 80.08 88.51 77.75 75.35 49.84 74.56 73.87
@@ F.3 Ablation Study
Figure 8: Additional changing trends of the averaged attention values with respect to training epochs.
F.4 Attention Trends.
For the analysis of attention trends, we provide visualizations on more datasets in Figure 8. It is noteworthy that, despite Pubmed being a homophilic graph with $\mathcal{H}=0.80$, non-local aggregation maintains a pivotal role in our SAF, diverging from the patterns observed on the Cora dataset as shown in Figure 2(c). This discrepancy ownes much to the sizable node count within Pubmed, where the capability of non-local aggregation in capturing long-range dependencies proves advantageous for improving model performance, as indicated in Table 2.
F.5 Time and Space Overheads
eigendecomposition. Our SAF framework pre-computes eigendecomposition once per graph and reuses it for efficient matrix inversion in Eq. (6). This aspect is crucial, as the forward-pass cost in model training often exceeds the preprocessing expense of eigendecomposition. To empirically validate this, we compare the time overheads of eigendecomposition with the training times of various models in Table 6. It is evident that for most datasets, the time consumed by decomposition is significantly less than the time required for model training. For larger graphs such as Pubmed and Amazon-ratings, although the decomposition time is greater than that for local GNNs like BernNet and ChebNetII, it is still faster than the more advanced global GNNs including SAF, NodeFormer, and GloGNN++. In case of space overheads, similar patterns can be observed.
Table 6: Time overheads (s) / Space overheads (MB). The second and third column blocks represent local and global GNNs, respectively.
@@ F.5 Time and Space Overheads
Method Cham. Squi. Texas Corn. Actor Cora Cite. Pubm. Amaz.
BernNet 8.36/72 13.74/232 3.92/5 4.16/5 4.88/292 5.24/64 5.52/152 6.06/1546 13.98/2389
ChebNetII 22.82/72 30.73/231 11.47/5 9.64/5 14.88/291 19.96/63 16.14/152 36.91/1584 22.50/2355
SAF 11.55/112 18.78/440 4.38/5 4.70/5 5.36/733 6.04/120 6.12/237 18.43/4515 47.47/6966
NodeFormer 58.96/1522 79.66/3965 14.29/15 18.89/37 66.20/775 19.25/480 32.00/764 68.57/2119 122.91/3056
GloGNN++ 35.63/290 68.31/1525 4.47/5 3.00/5 73.13/2471 32.68/331 12.35/607 5266.53/17892 3614.37/25260
Decomposition 0.58/141 1.59/540 0.02/1 0.02/1 3.93 /1206 1.00/140 0.77 /239 21.34/7641 40.88/11442
Table 7: Average running time per epoch (ms)/average total running time (s)
@@
Model Comparison. To evaluate the efficiency of our model, we compare the running times of our proposed variants, SAF and SAF-$\epsilon$, against four notable spectral GNNs (BernNet, JacobiConv, ChebNetII, Specformer) and two cutting-edge non-local GNNs (NodeFormer, GloGNN++), as detailed in Table 7. One can see that SAF, while slightly slower than its base model, BernNet, due to the integration of non-local modules, remains more efficient than other leading non-local GNNs. Among other top spectral GNNs, SAF manages to achieve a comparable efficiency. On the other hand, SAF-$\epsilon$ costs much larger time, a consequence of its quadratic complexity from sparsification. However, this trade-off allows SAF-$\epsilon$ to produce an high-quality new graph, enhancing the efficiency of non-local aggregation and better capturing long-range interconnections and graph heterophily.
Table 8: Full-supervised node classification accuracies (%). SAF-Bern and SAF-Cheb refers to SAF implementation respectively using BernNet and ChebNetII as base model.
Method Cham. Squi. Texas Corn. Actor Cora Cite. Pubm.
SAF-Bern 75.30$\pm$0.96 63.63$\pm$0.81 94.10$\pm$1.48 92.95$\pm$1.97 42.93$\pm$0.79 89.80$\pm$0.69 80.61$\pm$0.81 91.49$\pm$0.29
Improv. 6.77% 12.24% 1.48% 0.82% 1.22% 1.29% 0.53% 2.98%
Improv. 3.60% 6.34% 1.15% 0.32% 0.90% 0.85% 0.15% 2.34%
Appendix G Analysis of SAF with ChebNetII as Base Model
To expand the versatility of our SAF framework, we introduced ChebNetII as an alternative base model, chosen for its adherence to the non-negative constraint critical in our model design. The rationale behind this choice is ChebNetII’s use of Chebyshev interpolation to approximate Chebyshev polynomials, where ensuring non-negativity is simply a matter of keeping its learnable parameters $\{\gamma_{j}\}_{j=0}^{K}$ non-negative. Our experiments, as shown in Table 8, confirm that SAF can significantly enhances ChebNetII’s performance, underscoring the framework’s flexibility with different spectral filters.
Interestingly, we observed that SAF, when based on Bernstein polynomials (SAF-Bern), tends to slightly outperform SAF with ChebNetII (SAF-Cheb) in most datasets. The margin of improvement is also more pronounced with SAF-Bern. This phenomenon could be attributed to the $g_{\phi}(\lambda)\leq 1$ constraint within SAF (refer to Section 5), necessitating the rescaling of filter functions by their maximum graph spectrum values. For Bernstein polynomials, this maximum is readily obtained as the largest polynomial coefficient ($\max\{\phi_{k}\}_{k=0}^{K}$, as per Proposition 3). However, for Chebyshev polynomials, the best theoretical upper bound is the sum of absolute coefficients, $\sum_{k=0}^{K}|\phi_{k}|$, which is comparatively less precise. This difference may impact the quality of graph construction and, subsequently, the model’s performance. Exploring these nuances will be a focal point of our future research.

### 08 Feb 2024 v3

@@ Appendix B Eigendecomposition
Eigendecomposition breaks down a matrix into its eigenvalues and eigenvectors, offering insights into matrix properties, especially for the graph Laplacian. Despite computational demands, this technique has attracted surging interest in the graph learning community due to its theoretical richness, and it can be practically expedited for larger graphs using Lanczos (Lanczos, 1950) and Sparse Generalized Eigenvalue (Cai et al., 2021) algorithms. Recent advancements also underscore its value in various applications such as graph positional encoding (Belkin and Niyogi, 2003; Dwivedi et al., 2020), spectral graph convolution (Liao et al., 2019), graph domain adaptation (You et al., 2022), and graph robustness (Chang et al., 2021). For example, Laplacian eigenvectors have been widely used in identifying global position of nodes in the graph (Wang et al., 2022), particularly in recent popular graph transformers (Kreuzer et al., 2021; Kim et al., 2022; Rampášek et al., 2022), enhancing their expressiveness. Innovations like SignNet, BasisNet (Lim et al., 2022), and Sign Equivariant (Lim et al., 2023) have further optimized the processing of these eigenvectors. When exploring the expressive power of GNN models, Specformer (Bo et al., 2023) employs eigendecomposition for learning set-to-set graph filters, TEDGCN (Yan et al., 2023) leverages it for adaptive weighting of eigengraphs, and FE-GNN (Sun et al., 2023) taps into singular value decomposition (SVD) for graph feature expansion. In line with these developments, our method, SAF, also utilizes eigendecomposition to explicitly create a new graph, enabling efficient non-local aggregation with signed weights to tackle long-range dependency and graph heterophily.
Appendix C Dataset Information
@@
We conduct experiments on 13 real-world datasets from various domains. The detailed statistics are summarized in Table 3. Alongside standard data attributes, we also provide the longest geodesic (shortest-path) distance between graph nodes for better illustrating the non-local property that we investigate in Figures 1 and 7. Moreover, we adopt three metrics - edge homophily (Zhu et al., 2020b) $\mathcal{H}$, class homophily (Lim et al., 2021) $\mathcal{H}_{\text{class}}$, and adjusted homophily (Platonov et al., 2023a) $\mathcal{H}_{\text{adjusted}}$ to assess the graph’s homophily ratio, which ranges from 0 (high heterophily) to 1 (high homophily). While the first is a commonly used index, the latter two, considering class variability and potential imbalance, have been recently introduced for more accurate estimation. For our main text analysis regarding the adapted new graph, we primarily rely on the edge homophily metric, defined as =|$\mathcal{H}=|\{(v_{i},v_{j})|(v_{i},v_{j})\in\mathcal{E}\wedge\mathbf{y}_{i}=\mathbf{y}_{j}\}|/|\mathcal{E}|$, given its simplicity and wide usage. In certain compact sections of this paper, we use four-letter abbreviations for dataset names.
Appendix D Experimental Details
@@
In this section, we provide experimental details for reproducibility. As He et al. (2022) have made a comprehensive evaluation and share the same experimental protocol with us, we directly leverage their results for models: MLP, GCN, APPNP, ARMA, GPR-GNN, BernNet, ChebNetII, GCNII and PDE-GCN on datasets including Cham., Squi., Texas, Corn., Actor, Cora, Cite., and Pubm.. For JacobiConv, LON-GNN, and OptBasisGNN, we also report their results from corresponding papers (Wang and Zhang, 2022; Tao et al., 2023; Guo and Wei, 2023b). Other experiments are performed on a machine equipped with an NVIDIA GeForce RTX 3090 (24GB) and an Intel(R) Xeon(R) Gold 5218R CPU @ 2.10GHz (20 cores).
D.1. Baseline Implementations
@@
In our experiments, we leverage the Pytorch Geometric library (Fey and Lenssen, 2019) implementations for GCN and APPNP. For MLP, we include a sequence of linear layers, each of which is followed by batch normalization, ReLU activation, and dropout. The number of MLP layers are tuned from 1 to 5. For the remaining baselines, we resort to their publicably released code. Besides, for the work (Li et al., 2022), only the most effective model variant, GloGNN++, is included in our experiments. Models like Geom-GCN (Pei et al., 2020) and Non-Local GNNs (Liu et al., 2021), surpassed by these SOTA methods, are excluded from our comparison.
D.2. Hyper-parameters Setting
@@
Figure 7. Distributions of original hop distance between adjacent nodes in the adapted new graph on additional seven datasets.
Appendix E More Experiments
E.1. Empirical Evidence of the New Graph
Additional empirical support for the newfound graph’s properties is presented in Figures 7 and 8.
Figure 8. Additional visualizations for illustrating the newfound graph’s ability in discerning label consistency.
E.2. Attention Trends
For the analysis of attention trends, we provide visualizations on more datasets in Figure 9. It is noteworthy that, despite Pubmed being a homophilic graph with $\mathcal{H}=0.80$, non-local aggregation maintains a pivotal role in our SAF, diverging from the patterns observed on the Cora dataset as shown in Figure 5(b). This discrepancy ownes much to the sizable node count within Pubmed, where the capability of non-local aggregation in capturing long-range dependencies proves advantageous for improving model performance.
@@ E.2. Attention Trends
Figure 9. Additional changing trends of the averaged attention values with respect to training epochs.
E.3. New Benchmarks for Graph Heterophily
For a more extensive evaluation across various domains, we also test SAF on five recently introduced datasets, including Mine., Tolo., Amaz., Roma., and Penn94 (Platonov et al., 2023b; Lim et al., 2021). In this context, we draw comparisons solely with MLP, GCN, APPNP, along with six GNN models that have previously shown promising results in prior tasks, namely GPR-GNN, BernNet, ChebNetII, JacobiConv, FE-GNN, and ClenshawGCN. Table 4 lists the average classification accuracies, obtained over random splits provided by (Platonov et al., 2023b; Lim et al., 2021), with a distribution of 50%/25%/25% for training/validation/testing. In summary, SAF achieves significant performance gains of 12.84% and 4.10% on Mine. and Tolo. datasets, respectively, while maintaining competitiveness on the other three.
Table 4. Evaluations on new heterophilic graph datasets
Method Mine. Tolo. Amaz. Roma. Penn94
MLP 50.61$\pm$0.87 74.58$\pm$0.69 45.50$\pm$0.38 66.11$\pm$0.33 74.58$\pm$0.37
GCN 72.25$\pm$0.60 76.56$\pm$0.85 48.06$\pm$0.39 53.49$\pm$0.33 82.47$\pm$0.27
APPNP 68.48$\pm$1.20 74.13$\pm$0.62 48.12$\pm$0.37 72.99$\pm$0.46 75.29$\pm$0.27
GPR-GNN 89.76$\pm$0.53 75.82$\pm$0.50 49.06$\pm$0.25 73.19$\pm$0.24 81.38$\pm$0.16
ChebNetII 83.62$\pm$1.51 78.95$\pm$0.49 49.76$\pm$0.36 74.52$\pm$0.54 83.12$\pm$0.22
JacobiConv 89.88$\pm$0.33 77.24$\pm$0.39 43.89$\pm$0.28 74.30$\pm$0.50 83.35$\pm$0.11
FE-GNN 84.68$\pm$0.36 79.31$\pm$0.59 49.46$\pm$0.29 74.50$\pm$0.30 82.30$\pm$0.54
ClenshawGCN 90.36$\pm$0.92 80.94$\pm$0.52 50.14$\pm$0.52 73.14$\pm$0.51 84.71$\pm$0.31
BernNet 77.75$\pm$0.61 75.35$\pm$0.63 49.84$\pm$0.52 74.56$\pm$0.74 82.47$\pm$0.21
SAF 90.54$\pm$0.30 79.38$\pm$0.58 50.49$\pm$0.28 74.87$\pm$0.22 83.86$\pm$0.26
Improv. 12.79% 4.03% 0.65% 0.31% 1.39%
Table 5. Full-supervised node classification accuracy (%). SAF-Cheb refers to SAF implementation using Chebyshev polynomials.
@@ E.3. New Benchmarks for Graph Heterophily
Method Cham. Squi. Texas Corn. Actor Cora Cite. Pubm.
SAF-Cheb-$\epsilon$ 75.25$\pm$0.96 64.42$\pm$0.82 94.26$\pm$1.64 93.12$\pm$1.64 42.79$\pm$1.04 89.61$\pm$0.71 81.08$\pm$0.68 91.73$\pm$0.18
Improv. 3.88% 6.70% 1.15% 0.82% 1.04% 0.90% 0.15% 2.80%
E.4. SAF with ChebNetII as Base Model
To expand the versatility of our SAF framework, we introduced ChebNetII (He et al., 2022) as an alternative base model, chosen for its adherence to the non-negative constraint, critical in our model design as stated in Section 5.1. The rationale behind this choice is ChebNetII’s use of Chebyshev interpolation for learning Chebyshev polynomials, where the constraint can be ensured by keeping its learnable parameters $\{\gamma_{j}\}_{j=0}^{K}$ non-negative. Our experiments, as shown in Table 5, confirm that SAF can significantly enhances ChebNetII’s performance, underscoring the framework’s flexibility with different spectral filters. Interestingly, we observed that SAF, utilizing Bernstein polynomials (SAF-Bern), slightly surpasses its performance with Chebyshev polynomials (SAF-Cheb) in most datasets. The margin of improvement over the base model is also more pronounced with SAF-Bern. This phenomenon could be attributed to the $g_{\phi}(\lambda)\leq 1$ constraint within SAF (refer to Section 5.1), necessitating the rescaling of filter functions by their maximum values. For Bernstein polynomials, this maximum is readily obtained as the largest polynomial coefficient $\max\{\phi_{k}\}_{k=0}^{K}$, as per Proposition 3. However, for Chebyshev polynomials, the best theoretical upper bound is the sum of absolute coefficients, $\sum_{k=0}^{K}|\phi_{k}|$, which is comparatively less precise. This difference may impact the quality of graph construction and, subsequently, the model’s performance. Exploring these nuances will be a focal point of our future research.
Table 6. Time overheads (s) / Space overheads (MB). For large-scale graph Penn94 (with 41,554 nodes and 1,362,229 edges), we only employ a partial eigendecomposition with 100 extremal eigenvalues to balance between model effectiveness and efficiency.
Method Cham. Squi. Texas Corn. Actor Cora Cite. Pubm. Penn94
BernNet 8.36/72 13.74/232 3.92/5 4.16/5 4.88/292 5.24/64 5.52/152 6.06/1546 24.05/1902
ChebNetII 22.82/72 30.73/231 11.47/5 9.64/5 14.88/291 19.96/63 16.14/152 36.91/1584 41.67/1850
ClenshawGCN 15.58/98 26.21/398 2.98/5 4.13/5 4.43/313 3.72/68 12.48/152 24.97/1833 191.30/3017
SAF 11.55/112 18.78/440 4.38/5 4.70/5 5.36/733 6.04/120 6.12/237 18.43/4515 23.49/8491
Decomposition 0.58/141 1.59/540 0.02/1 0.02/1 3.93 /1206 1.00/140 0.77 /239 21.34/7641 4.76/4
Table 7. Average running time per epoch (ms) / average total running time (s).
@@ E.4. SAF with ChebNetII as Base Model
Table 8. Ablation study of SAF framework.
Our Variant Cham. Squi. Texas Corn. Actor Cora Cite. Pubm. Mine. Tolo. Amaz. Roma.
SAF-$\epsilon$ 74.84 64.00 94.75 93.28 42.98 89.87 81.45 91.52 90.59 79.45 50.36 74.89
SAF 75.30 63.63 94.10 92.95 42.93 89.80 80.61 91.49 90.54 79.38 50.49 74.87
SAF w/o Atte. 75.01 62.62 89.18 86.07 41.53 88.80 80.37 91.24 89.45 76.18 49.98 74.25
SAF w/o Spec. 73.55 55.70 90.49 88.20 41.06 88.03 78.87 90.12 89.45 78.23 49.13 71.85
SAF w/o Spat. 68.53 51.39 92.62 92.13 41.71 88.51 80.08 88.51 77.75 75.35 49.84 74.56
E.5. Time and Space Overheads
Eigendecomposition. Our SAF framework pre-computes eigendecomposition once per graph and reuses it in Eq. (6). This aspect is crucial, as the forward-pass cost in model training often exceeds the preprocessing expense of eigendecomposition. To empirically validate this, we compare the time overheads of eigendecomposition with the training times of various models in Table 6. It is evident that for most datasets, the time consumed by decomposition is significantly less than the time required for model training. For medium-sized graphs such as Pubmed, while the full decomposition time exceeds that of BernNet, it still maintains efficiency against more advanced GNNs (ChebNetII, ClenshawGCN, SAF). Moving to the large-scale graph, Penn94, where only partial eigendecomposition with 100 extremal eigenvalues is considered, the computation time is markedly reduced compared to all models mentioned.
Model Comparison. To evaluate the efficiency of our model, we compare the running times of our proposed variants, SAF and SAF-$\epsilon$, against four notable spectral GNNs (BernNet, ChebNetII, JacobiConv), two non-local GNNs (NodeFormer, GloGNN++), and one unified GNN model (ClenshawGCN), as detailed in Table 7. One can observe that SAF, while slightly slower than its base model, BernNet, due to the integration of non-local spatial aggregation, remains more efficient than or comparable to other SOTA methods. On the other hand, SAF-$\epsilon$ costs more time (but still faster than the non-local GNN, GloGNN++), a consequence of its quadratic complexity from non-local sparsification. However, this trade-off allows SAF-$\epsilon$ to produce an high-quality new graph, better capturing long-range interconnections among nodes and addressing graph heterophily.